# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

### v1

@@ 4.2 Pseudodifferential Operators and Energy Asymptotic
Fix $s\in\mathbb{R}$ and $\lambda>4\rho^{2}$ sufficiently large. Define the fractional Bessel potential space by
@@
We recall that $P_{s}$ with $s\in(0,\frac{n}{2})\setminus{\mathbb{N}}$ on $\mathbb{H}^{n}$ defined by spectral calculus as
$\mathcal{P}_{s}=m_{s}\!\bigl(\sqrt{-\mathcal{L}}\bigr),\qquad\mathcal{L}:=\Delta_{\mathbb{H}^{n}}+\rho^{2},$ (4.8)
where $\operatorname{Spec}(-\mathcal{L})=[0,\infty)$ and, in the Helgason–Fourier representation,
@@
Let $\gamma>0$ and set
$a:=\frac{3+2\gamma}{4}>0,\qquad b:=\frac{3-2\gamma}{4}.$
Define, for $\zeta\in\mathbb{C}$,
$m_{\gamma}(\zeta):=2^{2\gamma}\,\frac{\Gamma\!\bigl(a+\frac{i}{2}\zeta\bigr)\,\Gamma\!\bigl(a-\frac{i}{2}\zeta\bigr)}{\Gamma\!\bigl(b+\frac{i}{2}\zeta\bigr)\,\Gamma\!\bigl(b-\frac{i}{2}\zeta\bigr)}.$ (4.10)
Then $m_{\gamma}$ is holomorphic in every strip $|\Im\zeta|<W$ with $W<\epsilon$, where one may take
@@
Moreover, for each $W<\epsilon$ and each $k\in\mathbb{N}_{0}$ there exists a constant $C_{k,\gamma,W}>0$ such that
$|\partial_{\zeta}^{k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,W}\,(1+|\zeta|)^{2\gamma-k},\qquad|\Im\zeta|<W.$ (4.11)
Proof: The only possible singularities of (4.10) come from the poles of $\Gamma\!\bigl(a\pm\frac{i}{2}\zeta\bigr)$, which occur precisely when $a\pm\frac{i}{2}\zeta\in\{0,-1,-2,\dots\}$, i.e.
$\zeta=\pm 2i(a+k),\qquad k\in\mathbb{N}_{0}.$
Hence $m_{\gamma}$ is holomorphic in the strip $|\Im\zeta|<2a=:\epsilon$. Poles of $\Gamma\!\bigl(b\pm\frac{i}{2}\zeta\bigr)$, if any, produce zeros of $m_{\gamma}$ and therefore do not affect holomorphy.
Fix $W<\epsilon$ and write
$m_{\gamma}(\zeta)=2^{2\gamma}\,R_{+}(\zeta)\,R_{-}(\zeta),\qquad R_{\pm}(\zeta):=\frac{\Gamma\!\bigl(a\pm\frac{i}{2}\zeta\bigr)}{\Gamma\!\bigl(b\pm\frac{i}{2}\zeta\bigr)}.$
Note that $a=b+\gamma$. For $|\Im\zeta|<W$ and $\pm$ fixed, set
$z_{\pm}:=b\pm\frac{i}{2}\zeta,\qquad\text{so that}\qquad R_{\pm}(\zeta)=\frac{\Gamma(z_{\pm}+\gamma)}{\Gamma(z_{\pm})}.$
@@
$|R_{\pm}(\zeta)|\leq C_{\gamma,W}\,(1+|\zeta|)^{\gamma},\qquad|\Im\zeta|<W,$
@@
$|m_{\gamma}(\zeta)|\leq C_{\gamma,W}\,(1+|\zeta|)^{2\gamma},\qquad|\Im\zeta|<W.$
We next estimate derivatives for $|\zeta|\to\infty$ inside the strip. Choose $R_{0}\geq 1$ so large that for all $\zeta$ with $|\Im\zeta|<W$ and $|\zeta|\geq R_{0}$, the points $z_{\pm}$ stay at a positive distance from the poles $\{0,-1,-2,\dots\}$ of the digamma and polygamma functions. On the region
$\Omega_{R_{0},W}:=\{\zeta\in\mathbb{C}:\ |\Im\zeta|<W,\ |\zeta|\geq R_{0}\},$
@@
$\partial_{\zeta}\log R_{\pm}(\zeta)=\pm\frac{i}{2}\Bigl(\psi(z_{\pm}+\gamma)-\psi(z_{\pm})\Bigr).$
@@
$\psi(z_{\pm}+\gamma)-\psi(z_{\pm})=O\!\left(\frac{1}{|z_{\pm}|}\right)=O\!\left(\frac{1}{|\zeta|}\right),\qquad|\zeta|\to\infty,\quad\zeta\in\Omega_{R_{0},W}.$
@@
$\Bigl|\partial_{\zeta}^{\,j}\bigl(\partial_{\zeta}\log R_{\pm}(\zeta)\bigr)\Bigr|\leq C_{j,\gamma,W}\,(1+|\zeta|)^{-j-1},\qquad\zeta\in\Omega_{R_{0},W}.$
@@
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,W}\,|R_{\pm}(\zeta)|\,(1+|\zeta|)^{-k}\leq C_{k,\gamma,W}\,(1+|\zeta|)^{\gamma-k},\qquad\zeta\in\Omega_{R_{0},W}.$
@@
$K_{R_{0},W}:=\{\zeta\in\mathbb{C}:\ |\Im\zeta|\leq W,\ |\zeta|\leq R_{0}\}.$
Since $m_{\gamma}$ (hence $R_{\pm}$) is holomorphic in a neighbourhood of $K_{R_{0},W}$, each derivative $\partial_{\zeta}^{\,k}R_{\pm}$ is continuous and bounded on $K_{R_{0},W}$. Thus there exists $C_{k,\gamma,W}>0$ such that
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,W}\,(1+|\zeta|)^{\gamma-k},\qquad\zeta\in K_{R_{0},W}.$
Combining this with the estimate on $\Omega_{R_{0},W}$ yields, for all $\zeta$ with $|\Im\zeta|<W$,
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,W}\,(1+|\zeta|)^{\gamma-k}.$
@@
$|\partial_{\zeta}^{\,k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,W}\sum_{j=0}^{k}(1+|\zeta|)^{\gamma-j}(1+|\zeta|)^{\gamma-(k-j)}\leq C_{k,\gamma,W}\,(1+|\zeta|)^{2\gamma-k},\qquad|\Im\zeta|<W,$
@@
The above conclusion also holds for the multiplier $\widetilde{m}_{\gamma}$ (see 2.8), the proof is entirely analogous and will be omitted.
@@
Let $\gamma\in\bigl(0,\frac{n}{2}\bigr)$. Then both operators $P_{\gamma}$ and $\widetilde{P}_{\gamma}$ extend by continuity to bounded linear maps
@@
$T_{\lambda}:=(\lambda I-\Delta)^{\frac{s-2\gamma}{2}}\,\mathcal{P}_{\gamma}\,(\lambda I-\Delta)^{-\frac{s}{2}}:L^{2}(\mathbb{H}^{n})\to L^{2}(\mathbb{H}^{n}),$ (4.13)
where $\Delta=\Delta_{\mathbb{H}^{n}}$. Indeed, if $u=(\lambda I-\Delta)^{-s/2}f$ with $f\in L^{2}$, then
$\|\mathcal{P}_{\gamma}u\|_{\tilde{H}^{s-2\gamma}}=\bigl\|(\lambda I-\Delta)^{\frac{s-2\gamma}{2}}P_{\gamma}(\lambda I-\Delta)^{-\frac{s}{2}}f\bigr\|_{L^{2}}=\|T_{\lambda}f\|_{L^{2}}.$
Recall $\mathcal{L}=\Delta+\rho^{2}$, hence
$\lambda I-\Delta=(\lambda+\rho^{2})I-\mathcal{L},$ (4.14)
see [30, Scetion 5]. Using (4.8) and (4.14), we may view every factor in (4.13) as a function of the same self-adjoint operator $\sqrt{-\mathcal{L}}$. In spectral variable $\zeta\geq 0$ (corresponding to $\sqrt{-\mathcal{L}}$), we obtain
$T_{\lambda}=\Psi_{\lambda}\!\bigl(\sqrt{-\mathcal{L}}\bigr),\qquad\Psi_{\lambda}(\zeta):=(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}\,m_{\gamma}(\zeta).$
Let $\mathcal{S}_{W}^{m}$ denote the standard strip-holomorphic symbol class appearing in the $L^{p}$ multiplier theorem on noncompact manifolds with bounded geometry: namely $\Phi\in\mathcal{S}_{W}^{m}$ if $\Phi$ is an even function and extends holomorphically to $\{\zeta\in\mathbb{C}:|\Im\zeta|<W\}$ and satisfies $|\partial_{\zeta}^{k}\Phi(\zeta)|\leq C_{k}(1+|\zeta|)^{m-k}$ there. By [30, Chapter IV, Proposition 2.4], it is enough to show that there exists $W>0$ such that
$\Psi_{\lambda}\in\mathcal{S}_{W}^{0}.$ (4.15)
Indeed, taking sufficiently small $\epsilon>0,$ for $W=\frac{\epsilon}{2}<\epsilon$, the functions $\zeta\mapsto(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}$ are holomorphic in $|\Im\zeta|<W$, and a direct differentiation shows
$\bigl|\partial_{\zeta}^{k}(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}\bigr|\leq C_{k,\gamma,\lambda,\rho,W}\,(1+|\zeta|)^{-2\gamma-k},\qquad|\Im\zeta|<W,$
@@
By Lemma 4.5, the multiplier $m_{\gamma}$ in (4.9) is even and extends holomorphically to every strip $|\Im\zeta|<W<\epsilon$ and satisfies
$|\partial_{\zeta}^{k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,\rho,W}\,(1+|\zeta|)^{2\gamma-k},\qquad|\Im\zeta|<W.$
Hence $m_{\gamma}\in S_{\frac{\epsilon}{2}}^{2\gamma}$. Multiplying the two symbols yields (4.15), see [16, Chapter 18]
$\Psi_{\lambda}(\zeta)=(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}m_{\gamma}(\zeta)\in S_{\frac{\epsilon}{4}}^{0}.$
This completes the proof for $P_{\gamma}$. The argument for $\widetilde{P}_{\gamma}$ is entirely analogous. ∎
@@
Let $s\in\bigl(0,\frac{n}{2}\bigr)$ and let $\phi\in C_{c}^{\infty}(\mathbb{H}^{n})$. Define the commutator
$\mathcal{P}_{s},\phi]:=\mathcal{P}_{s}\circ M_{\phi}-M_{\phi}\circ\mathcal{P}_{s},\qquad M_{\phi}u:=\phi u.$
Then $\mathcal{P}_{s},\phi]$ extends by continuity to a bounded operator
$\mathcal{P}_{s},\phi]:\widetilde{H}^{\,s-1}(\mathbb{H}^{n})\longrightarrow\widetilde{H}^{-s}(\mathbb{H}^{n}),\qquad\|[\mathcal{P}_{s},\phi]f\|_{\widetilde{H}^{-s}}\leq C_{\phi}\,\|f\|_{\widetilde{H}^{\,s-1}},$ (4.16)
for some constant $C_{\phi}>0$. Moreover, the same conclusion holds with $\mathcal{P}_{s}$ replaced by $\widetilde{\mathcal{P}}_{s}$.
Proof: From the proof of Proposition 4.6, for any $W>0$ small enough one has
$m_{s}\in S_{W}^{2s}.$
By [30, Chapter IV, Proposition 1.2], it is equivalent to $\mathcal{P}_{s}\in\Psi_{W}^{2s}(\mathbb{H}^{n}).$
Multiplication by $\phi$ is a zero-order operator, $M_{\phi}\in\Psi_{W}^{0}(\mathbb{H}^{n})$. By the standard symbolic calculus for pseudodifferential operators, the commutator satisfies
$\mathcal{P}_{s},M_{\phi}]\in\Psi_{W}^{2s-1}(\mathbb{H}^{n}).$
@@
$\sigma([\mathcal{P}_{s},M_{\phi}])(x,\xi)\sim\sum_{|\alpha|\geq 1}\frac{1}{\alpha!}\,\partial_{\xi}^{\alpha}p_{s}(x,\xi)\,D_{x}^{\alpha}\phi(x),$
@@
Finally, we use the Sobolev mapping property for pseudodifferential operators ([30, Proposition 4.5]): for $T\in\Psi_{W}^{m}(\mathbb{H}^{n})$ and any $t\in\mathbb{R}$,
$T:\tilde{H}^{t}(\mathbb{H}^{n})\longrightarrow\tilde{H}^{t-m}(\mathbb{H}^{n})\quad\text{bounded}.$ (4.17)
Taking $T=[\mathcal{P}_{s},\phi]$, m=2s1m=2s-1 and t=s1t=s-1 in (4.17), we obtain
$\mathcal{P}_{s},\phi]:\tilde{H}^{s-1}(\mathbb{H}^{n})\to\tilde{H}^{-s}(\mathbb{H}^{n}),$
together with the estimate (4.16). The constant $C_{\phi}$ depends only on $\phi$. ∎
@@
Let $\{v_{j}\}\subset C_{c}^{\infty}(\mathbb{H}^{n})$ be a bounded sequence and assume that $v_{j}\rightharpoonup 0$ weakly in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$. Then, for every $\phi\in C_{c}^{\infty}(\mathbb{H}^{n})$, the following hold as $j\to\infty$:
@@
    $\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}=\mathcal{E}_{\lambda,s}(\phi v_{j})+o(1).$
@@
    $\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s,\sim}=\widetilde{\mathcal{E}}_{\lambda,s}(\phi v_{j})+o(1).$

### v2

@@ 4.2 Pseudodifferential Operators and Energy Asymptotic
In this subsection, we establish the pseudodifferential tools needed for the energy asymptotic analysis. We first study the multiplier $m_{\gamma}$, proving strip holomorphy and symbol-type derivative bounds, which place $\mathcal{P}_{\gamma}$ and $\widetilde{\mathcal{P}}_{\gamma}$ within a standard pseudodifferential framework. As a consequence, we obtain Sobolev mapping properties between spectral spaces of different orders. We then derive commutator estimates with compactly supported cutoffs, showing that these commutators are bounded lower-order operators. Finally, combined with weak convergence in the energy space, these estimates yield an asymptotic localization identity: two natural cut-off localizations of the same nonlocal quadratic form are equivalent up to $o(1)$-errors. These tools form the technical basis for the concentration analysis and the variational arguments in the sequel.
We start from recalling the fractional Bessel potential spaces, see [42]. Fix $s\in\mathbb{R}$ and $\lambda>4\rho^{2}$ sufficiently large. Define the fractional Bessel potential space by
@@
We recall that $\mathcal{P}_{s}$ with $s\in(0,\frac{n}{2})\setminus{\mathbb{N}}$ on $\mathbb{H}^{n}$ defined by spectral calculus as
$\mathcal{P}_{s}=m_{s}\!\bigl(\mathcal{A}\bigr)\quad{\rm with}\ \,\mathcal{A}=\sqrt{-\Delta_{\mathbb{H}^{n}}-\rho^{2}}$ (4.8)
where $\operatorname{Spec}(\mathcal{A}^{2})=[0,\infty)$ and, in the Helgason–Fourier representation,
@@
Let
$\gamma>0,\qquad a=\frac{3+2\gamma}{4}>0,\qquad b=\frac{3-2\gamma}{4}$
and
$m_{\gamma}(\zeta):=2^{2\gamma}\,\frac{\Gamma\!\bigl(a+\frac{{\bf i}}{2}\zeta\bigr)\,\Gamma\!\bigl(a-\frac{{\bf i}}{2}\zeta\bigr)}{\Gamma\!\bigl(b+\frac{{\bf i}}{2}\zeta\bigr)\,\Gamma\!\bigl(b-\frac{{\bf i}}{2}\zeta\bigr)}\quad\text{for $\zeta\in\mathbb{C}$}.$ (4.10)
Then $m_{\gamma}$ is holomorphic in every strip $|{\rm Im}(\zeta)|<\sigma$ with $\sigma<\epsilon$, where one may take
@@
Moreover, for each $\sigma<\epsilon$ and each $k\in\mathbb{N}_{0}$ there exists a constant $C_{k,\gamma,\sigma}>0$ such that
$|\partial_{\zeta}^{k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,\sigma}\,(1+|\zeta|)^{2\gamma-k},\qquad|{\rm Im}(\zeta)|<\sigma,$ (4.11)
where ${\rm Im}(\zeta)=\zeta_{2}$ if $\zeta=\zeta_{1}+{\bf i}\zeta_{2}$, $\zeta_{1},\zeta_{2}\in\mathbb{R}$.
Proof: The only possible singularities of (4.10) come from the poles of $\Gamma\!\bigl(a\pm\frac{{\bf i}}{2}\zeta\bigr)$, which occur precisely when $a\pm\frac{{\bf i}}{2}\zeta\in\{0,-1,-2,\dots\}$, i.e.
$\zeta=\pm 2{\bf i}(a+k),\qquad k\in\mathbb{N}_{0}.$
Hence $m_{\gamma}$ is holomorphic in the strip $|{\rm Im}(\zeta)|<2a=:\epsilon$. Poles of $\Gamma\!\bigl(b\pm\frac{{\bf i}}{2}\zeta\bigr)$, if any, produce zeros of $m_{\gamma}$ and therefore do not affect holomorphy.
Fix $\sigma<\epsilon$ and write
$m_{\gamma}(\zeta)=2^{2\gamma}\,R_{+}(\zeta)\,R_{-}(\zeta),\qquad R_{\pm}(\zeta):=\frac{\Gamma\!\bigl(a\pm\frac{{\bf i}}{2}\zeta\bigr)}{\Gamma\!\bigl(b\pm\frac{{\bf i}}{2}\zeta\bigr)}.$
Note that $a=b+\gamma$. For $|{\rm Im}(\zeta)|<\sigma$ and $\pm$ fixed, set
$z_{\pm}:=b\pm\frac{{\bf i}}{2}\zeta,\qquad\text{so that}\qquad R_{\pm}(\zeta)=\frac{\Gamma(z_{\pm}+\gamma)}{\Gamma(z_{\pm})}.$
@@
$|R_{\pm}(\zeta)|\leq C_{\gamma,\sigma}\,(1+|\zeta|)^{\gamma},\qquad|{\rm Im}(\zeta)|<\sigma,$
@@
$|m_{\gamma}(\zeta)|\leq C_{\gamma,\sigma}\,(1+|\zeta|)^{2\gamma},\qquad|{\rm Im}(\zeta)|<\sigma.$
We next estimate derivatives for $|\zeta|\to\infty$ inside the strip. Choose $R_{0}\geq 1$ so large that for all $\zeta$ with $|{\rm Im}(\zeta)|<\sigma$ and $|\zeta|\geq R_{0}$, the points $z_{\pm}$ stay at a positive distance from the poles $\{0,-1,-2,\dots\}$ of the digamma and polygamma functions. On the region
$\Omega_{R_{0},\sigma}:=\{\zeta\in\mathbb{C}:\ |{\rm Im}(\zeta)|<\sigma,\ |\zeta|\geq R_{0}\},$
@@
$\partial_{\zeta}\log R_{\pm}(\zeta)=\pm\frac{{\bf i}}{2}\Bigl(\psi(z_{\pm}+\gamma)-\psi(z_{\pm})\Bigr).$
@@
$\psi(z_{\pm}+\gamma)-\psi(z_{\pm})=O\!\left(\frac{1}{|z_{\pm}|}\right)=O\!\left(\frac{1}{|\zeta|}\right)\quad{\rm for}\ \,|\zeta|\to\infty,\quad\zeta\in\Omega_{R_{0},W}.$
@@
$\Bigl|\partial_{\zeta}^{\,j}\bigl(\partial_{\zeta}\log R_{\pm}(\zeta)\bigr)\Bigr|\leq C_{j,\gamma,\sigma}\,(1+|\zeta|)^{-j-1},\qquad\zeta\in\Omega_{R_{0},\sigma}.$
@@
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,\sigma}\,|R_{\pm}(\zeta)|\,(1+|\zeta|)^{-k}\leq C_{k,\gamma,\sigma}\,(1+|\zeta|)^{\gamma-k},\qquad\zeta\in\Omega_{R_{0},\sigma}.$
@@
$K_{R_{0},\sigma}:=\big\{\zeta\in\mathbb{C}:\ |{\rm Im}(\zeta)|\leq\sigma,\ |\zeta|\leq R_{0}\big\}.$
Since $m_{\gamma}$ (hence $R_{\pm}$) is holomorphic in a neighbourhood of $K_{R_{0},W}$, each derivative $\partial_{\zeta}^{\,k}R_{\pm}$ is continuous and bounded on $K_{R_{0},\sigma}$. Thus there exists $C_{k,\gamma,\sigma}>0$ such that
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,\sigma}\,(1+|\zeta|)^{\gamma-k},\qquad\zeta\in K_{R_{0},\sigma}.$
Combining this with the estimate on $\Omega_{R_{0},\sigma}$ yields, for all $\zeta$ with $|{\rm Im}(\zeta)|<\sigma$,
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,\sigma}\,(1+|\zeta|)^{\gamma-k}.$
@@
$|\partial_{\zeta}^{\,k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,W}\sum_{j=0}^{k}(1+|\zeta|)^{\gamma-j}(1+|\zeta|)^{\gamma-(k-j)}\leq C_{k,\gamma,\sigma}\,(1+|\zeta|)^{2\gamma-k},\qquad|{\rm Im}(\zeta)|<\sigma,$
@@
The above conclusion also holds for the multiplier $\widetilde{m}_{\gamma}$ (see 2.9), the proof is entirely analogous and will be omitted.
The symbol estimates obtained above immediately yield the Sobolev mapping properties of the fractional operators. In particular, both $\mathcal{P}_{\gamma}$ and $\widetilde{\mathcal{P}}_{\gamma}$ act as operators of order $2\gamma$ on the hyperbolic Bessel potential scale.
@@
Let $\gamma\in\bigl(0,\frac{n}{2}\bigr)$, then both operators $\mathcal{P}_{\gamma}$ and $\widetilde{\mathcal{P}}_{\gamma}$ extend by continuity to bounded linear maps
@@
${\bf T}_{\lambda}:=(\lambda I-\Delta_{\mathbb{H}^{n}})^{\frac{s-2\gamma}{2}}\,\mathcal{P}_{\gamma}\,(\lambda I-\Delta_{\mathbb{H}^{n}})^{-\frac{s}{2}}:L^{2}(\mathbb{H}^{n})\to L^{2}(\mathbb{H}^{n}).$ (4.13)
Indeed, if $u=(\lambda I-\Delta_{\mathbb{H}^{n}})^{-s/2}f$ with $f\in L^{2}$, then
$\|\mathcal{P}_{\gamma}u\|_{\tilde{H}^{s-2\gamma}}=\bigl\|(\lambda I-\Delta_{\mathbb{H}^{n}})^{\frac{s-2\gamma}{2}}\mathcal{P}_{\gamma}(\lambda I-\Delta_{\mathbb{H}^{n}})^{-\frac{s}{2}}f\bigr\|_{L^{2}}=\|{\bf T}_{\lambda}f\|_{L^{2}}.$
Note that
$\lambda I-\Delta_{\mathbb{H}^{n}}=(\lambda+\rho^{2})I+\mathcal{A}^{2},$ (4.14)
see [42, Scetion 5]. Using (4.8) and (4.14), we may view every factor in (4.13) as a function of the same self-adjoint operator $\sqrt{-\mathcal{L}}$. In spectral variable $\zeta\geq 0$ (corresponding to $\sqrt{-\mathcal{L}}$), we obtain
${\bf T}_{\lambda}=\Psi_{\lambda}\!\bigl(\mathcal{A}\bigr),\qquad\Psi_{\lambda}(\zeta):=(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}\,m_{\gamma}(\zeta).$
Let $\mathcal{S}_{\sigma}^{m}$ denote the standard strip-holomorphic symbol class appearing in the $L^{p}$ multiplier theorem on noncompact manifolds with bounded geometry: namely $\Phi\in\mathcal{S}_{\sigma}^{m}$ if $\Phi$ is an even function and extends holomorphically to $\{\zeta\in\mathbb{C}:|{\rm Im}(\zeta)|<\sigma\}$ and satisfies $|\partial_{\zeta}^{k}\Phi(\zeta)|\leq C_{k}(1+|\zeta|)^{m-k}$ there. By [42, Chapter IV, Proposition 2.4], it is enough to show that there exists $\sigma>0$ such that
$\Psi_{\lambda}\in\mathcal{S}_{\sigma}^{0}.$ (4.15)
Indeed, taking sufficiently small $\epsilon>0,$ for $\sigma=\frac{\epsilon}{2}<\epsilon$, the functions $\zeta\mapsto(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}$ are holomorphic in $|{\rm Im}(\zeta)|<W$, and a direct differentiation shows
$\bigl|\partial_{\zeta}^{k}(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}\bigr|\leq C_{k,\gamma,\lambda,\rho,\sigma}\,(1+|\zeta|)^{-2\gamma-k},\qquad|{\rm Im}(\zeta)|<\sigma,$
@@
By Lemma 4.5, the multiplier $m_{\gamma}$ in (4.9) is even and extends holomorphically to every strip $|{\rm Im}(\zeta)|<\sigma<\epsilon$ and satisfies
$|\partial_{\zeta}^{k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,\rho,\sigma}\,(1+|\zeta|)^{2\gamma-k},\qquad|{\rm Im}(\zeta)|<\sigma.$
Hence $m_{\gamma}\in\mathcal{S}_{\frac{\epsilon}{2}}^{2\gamma}$. Multiplying the two symbols yields (4.15), see [20, Chapter 18]
$\Psi_{\lambda}(\zeta)=(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}m_{\gamma}(\zeta)\in\mathcal{S}_{\frac{\epsilon}{4}}^{0}.$
This completes the proof for $\mathcal{P}_{\gamma}$. The argument for $\widetilde{\mathcal{P}}_{\gamma}$ is entirely analogous. ∎
The next result provides the required commutator estimate: the commutator with a smooth compactly supported multiplier is of lower order and is bounded between the corresponding hyperbolic Bessel potential spaces.
@@
Let $s\in\bigl(0,\frac{n}{2}\bigr)$ and the commutator
$\mathcal{P}_{s},\psi]:=\mathcal{P}_{s}\circ M_{\psi}-M_{\psi}\circ\mathcal{P}_{s}\quad\,{\rm with}\ \,M_{\psi}u:=\psi u$
for $\psi\in C_{c}^{\infty}(\mathbb{H}^{n})$. Then $\mathcal{P}_{s},\psi]$ extends by continuity to a bounded operator
$\mathcal{P}_{s},\psi]:\widetilde{H}^{\,s-1}(\mathbb{H}^{n})\longrightarrow\widetilde{H}^{-s}(\mathbb{H}^{n}),\qquad\|[\mathcal{P}_{s},\psi]f\|_{\widetilde{H}^{-s}}\leq C_{\psi}\,\|f\|_{\widetilde{H}^{\,s-1}},$ (4.16)
for some constant $C_{\psi}>0$. Moreover, the same conclusion holds with $\mathcal{P}_{s}$ replaced by $\widetilde{\mathcal{P}}_{s}$.
Proof: From the proof of Proposition 4.6, for any $\sigma>0$ small enough one has
$m_{s}\in\mathcal{S}_{\sigma}^{2s}.$
By [42, Chapter IV, Proposition 1.2], it is equivalent to $\mathcal{P}_{s}\in\Psi_{\sigma}^{2s}(\mathbb{H}^{n}),$ where the space $\Psi_{W}^{m}(\mathbb{H}^{n})$, of a class of operators whose Schwartz kernels behave like those of operators in $\mathcal{S}^{m}_{0,1}$ near the diagonal, in a uniform fashion, and away from the diagonal decay like $d(x,y)^{-k}e^{-Wd(x,y)},\ \forall k$ as do all derivatives. See [42, Page 72].
Multiplication by $\psi$ is a zero-order operator, $M_{\psi}\in\Psi_{\sigma}^{0}(\mathbb{H}^{n})$. By the standard symbolic calculus for pseudodifferential operators, the commutator satisfies
$\mathcal{P}_{s},M_{\psi}]\in\Psi_{\sigma}^{2s-1}(\mathbb{H}^{n}).$
@@
$\sigma([\mathcal{P}_{s},M_{\psi}])(x,\xi)\sim\sum_{|\alpha|\geq 1}\frac{1}{\alpha!}\,\partial_{\xi}^{\alpha}p_{s}(x,\xi)\,D_{x}^{\alpha}\psi(x),$
@@
Finally, we use the Sobolev mapping property for pseudo-differential operators ([42, Proposition 4.5]): for ${\bf T}\in\Psi_{\sigma}^{m}(\mathbb{H}^{n})$ and any $t\in\mathbb{R}$,
${\bf T}:\tilde{H}^{t}(\mathbb{H}^{n})\longrightarrow\tilde{H}^{t-m}(\mathbb{H}^{n})\quad\text{bounded}.$ (4.17)
Taking ${\bf T}=[\mathcal{P}_{s},\psi]$, m=2s1m=2s-1 and t=s1t=s-1 in (4.17), we obtain
$\mathcal{P}_{s},\psi]:\tilde{H}^{s-1}(\mathbb{H}^{n})\to\tilde{H}^{-s}(\mathbb{H}^{n}),$
together with the estimate (4.16). The constant $C_{\psi}$ depends only on $\psi$. ∎
We now establish an asymptotic localization identity for weakly convergent sequences. Combining the commutator estimate with boundedness in the energy space, we show that for a compactly supported cut-off, the two natural localizations of the same nonlocal quadratic form are equivalent up to an $o(1)$-error. This fact will be used in the concentration–compactness analysis.
@@
Let $\{v_{j}\}\subset C_{c}^{\infty}(\mathbb{H}^{n})$ be a bounded sequence and assume that $v_{j}\rightharpoonup 0$ weakly in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$. Then, for every $\psi\in C_{c}^{\infty}(\mathbb{H}^{n})$, the following hold as $j\to\infty$:
@@
    $\langle v_{j},\psi^{2}v_{j}\rangle_{\lambda,s}=\mathcal{E}_{\lambda,s}(\psi v_{j})+o(1).$
@@
    $\langle v_{j},\psi^{2}v_{j}\rangle_{\lambda,s,\sim}=\widetilde{\mathcal{E}}_{\lambda,s}(\psi v_{j})+o(1).$